# Correlated decoding of logical algorithms with transversal gates

### v1

@@ 2 Numerical simulations
The logical error rate is estimated by Monte-Carlo sampling the physical errors over the chosen noise model, then generating the syndrome and logical observable measurements corresponding to the sampled set of physical errors. The decoders then predict which errors occurred based on the syndrome measurements, and correct the logical observable based on this prediction. A logical error occurs if the corrected logical observable measurement differs from the true, noiseless value. In this work, we only simulate Clifford circuits with a noise model consisting of Pauli errors, with the exception of Section 2.4. These simulations can be performed in polynomial time in the number of physical qubits, according to the Gottesman-Knill theorem Gottesman (1997). We use the open-source Clifford simulation package, Stim Gidney (2021), to simulate the circuits studied in this work. Unless otherwise stated, the error bars for the numerically computed logical error rates are the Clopper-Pearson confidence interval based on a Beta distribution with a significance level of $0.05$.
@@ 2.1 Surface code Bell state with perfect syndrome measurements
As with the other logical circuits in this work, we use Stim to construct the decoding hypergraph. Because the codes are initialized perfectly, we define the checks simply as the stabilizers measured at the end of the circuit, as no information is gained from comparing the final stabilizer measurement with the $+1$ stabilizers during state preparation. These checks form the vertices of the hypergraph. The hyperedges correspond to the possible single-physical-qubit $X$ and $Z$ errors occuring before or after the CNOT. As illustrated in Fig. 1(b) of the main text, physical $X$ $(Z)$ errors before the CNOT on the control (target) qubit are transferred to the other logical qubit, in the bulk flipping four checks with probability $pf_{b}$, two for each code. Physical $X$ $(Z)$ errors before the CNOT on the target (control) qubit are not transferred to the other qubit, as well as errors after the CNOT. In the bulk, these errors only flip two checks on one surface code. Because MWPM cannot handle error sources that flip more than two checks, we decompose the hyperedges that flip four checks into two existing hyperedges that flip two checks within a code, using the $\mathtt{decompose\_errors=True}$ hyperparameter in Stim. Because these edges already exist in the hypergraph, the error probabilities of the decomposed edges are combined with the existing error probabilities (see Ref. Gidney (2021)).
@@ 2.2 Surface code Bell state with $2d$ rounds of noisy syndrome measurements
Throughout the circuit, we apply the circuit-level noise model in Ref. Higgott et al. (2023) to each physical operation. In this noise model, each physical two-qubit gate is followed by uniform two-qubit depolarizing noise with probability $p$. Qubits involved in a single-qubit gate or qubits which idle during a gate duration experience a single-physical-qubit depolarizing channel with probability $p$. Physical qubits initialized in the $X$ ($Z$) basis are followed by a $Z$ ($X$) error with probability $32p/$ 3, and measurements in the $X$ ($Z$) basis are preceded by a $Z$ ($X$) error with probability $32p/$ 3 .
@@ 2.3 Deep logical transversal Clifford circuits
$C$
1.13(21)
0.464(65)
$\alpha_{\frac{1}{8}}$
0.097(7)
$\alpha_{\frac{1}{4}}$
0.070(5)
0.458(19)
$\alpha_{\frac{1}{2}}$
0.055(4)
0.240(10)
$\alpha_{1}$
0.049(3)
0.170(7)
$\alpha_{2}$
0.039(3)
0.146(6)
$\alpha_{3}$
0.039(3)
0.120(5)
$\alpha_{4}$
0.044(2)
0.109(4)
$\alpha_{5}$
0.041(2)
0.102(4)
$\alpha_{6}$
0.103(4)
$\alpha_{7}$
0.097(4)

### v2

@@ 2 Numerical simulations
The logical error rate is estimated by Monte-Carlo sampling the physical errors over the chosen noise model, then generating the syndrome and logical observable measurements corresponding to the sampled set of physical errors. The decoders then predict which errors occurred based on the syndrome measurements, and correct the logical observable based on this prediction. A logical error occurs if the corrected logical observable measurement differs from the true, noiseless value. In this work, we only simulate Clifford circuits with a noise model consisting of Pauli errors, with the exception of Section 2.4. These simulations can be performed in polynomial time in the number of physical qubits, according to the Gottesman-Knill theorem Gottesman (1997). We use the open-source Clifford simulation package, Stim Gidney (2021), to simulate the circuits studied in this work. Unless otherwise stated, the error bars for the numerically computed logical error rates are the Clopper-Pearson confidence interval based on a Beta distribution with a significance level of $0.05$. This choice corresponds to a 95% likelihood that the true logical error rate lies between the error bars, and is comparable to significance levels used in prior works Wu et al. (2022); Bausch et al. (2023).
@@ 2.1 Surface code Bell state with perfect syndrome measurements
As with the other logical circuits in this work, we use Stim to construct the decoding hypergraph. Because the codes are initialized perfectly, we define the checks simply as the stabilizers measured at the end of the circuit, as no information is gained from comparing the final stabilizer measurement with the $+1$ stabilizers during state preparation. These checks form the vertices of the hypergraph. The hyperedges correspond to the possible single-physical-qubit $X$ and $Z$ errors occuring before or after the CNOT. As illustrated in Fig. 1(b) of the main text, physical $X$ $(Z)$ errors before the CNOT on the control (target) qubit are transferred to the other logical qubit, in the bulk flipping four checks with probability $pf_{b}$, two for each code. Physical $X$ $(Z)$ errors before the CNOT on the target (control) qubit are not transferred to the other qubit, as well as errors after the CNOT. In the bulk, these errors only flip two checks on one surface code. Because MWPM cannot handle error sources that flip more than two checks, we decompose the hyperedges that flip four checks into two existing hyperedges that flip two checks within a code. We perform the decomposition by setting decompose_errors = True in Stim. This setting attempts to decompose each hyperedge that flips more than two checks into existing edges which flip one or two checks, if possible. Because these decomposed edges already exist in the hypergraph, the error probabilities of the decomposed edges are combined with the existing error probabilities (see Ref. Gidney (2021)).
@@ 2.2 Surface code Bell state with $2d$ rounds of noisy syndrome measurements
Throughout the circuit, we apply the circuit-level noise model in Ref. Higgott et al. (2023) to each physical operation. This same noise model is considered in the deep logical transversal Clifford circuits in Section 2.3. Although this noise model is similar to standard noise models considered in prior theoretical results, we note that these decoding techniques have also been applied to realistic device noise models in the experiments of Ref. Bluvstein et al. (2024). In this noise model, each physical two-qubit gate is followed by uniform two-qubit depolarizing noise with probability $p$. Qubits involved in a single-qubit gate or qubits which idle during a gate duration experience a single-physical-qubit depolarizing channel with probability $p$. Physical qubits initialized in the $X$ ($Z$) basis are followed by a $Z$ ($X$) error with probability $32p/$ 3, and measurements in the $X$ ($Z$) basis are preceded by a $Z$ ($X$) error with probability $32p/$ 3 .
@@ 2.3 Deep logical transversal Clifford circuits
$C$ 1.13(21) 0.464(65)
$\alpha_{\frac{1}{8}}$ 0.097(7)
$\alpha_{\frac{1}{4}}$ 0.070(5) 0.458(19)
$\alpha_{\frac{1}{2}}$ 0.055(4) 0.240(10)
$\alpha_{1}$ 0.049(3) 0.170(7)
$\alpha_{2}$ 0.039(3) 0.146(6)
$\alpha_{3}$ 0.039(3) 0.120(5)
$\alpha_{4}$ 0.044(2) 0.109(4)
$\alpha_{5}$ 0.041(2) 0.102(4)
$\alpha_{6}$ 0.103(4)
$\alpha_{7}$ 0.097(4)